OpenStudy (anonymous):

help and i will give a medal In ΔABC, if the lengths of sides a and c are 8 centimeters and 16 centimeters, respectively, and the measure of ∠C is 35°, what is the measure of ∠A to two decimal places? a.16.67° b.17.50° c.26.57° d.35.53

OpenStudy (trojanpoem):



OpenStudy (anonymous):

what forula should i use sine or cosine

OpenStudy (anonymous):

forula*

OpenStudy (anonymous):

formula*

OpenStudy (trojanpoem):

\[\frac{ a }{ \sin a } = \frac{ c }{ \sin c }\] \[\sin a = \frac{ a \sin c }{ c } = \frac{ 8 \times \sin 35 }{ 16 } \approx 0.287\] \[\sin^{-1} a = 16, 39\]

OpenStudy (anonymous):

so the answer would be 16.67

OpenStudy (trojanpoem):

Sine will be easier although no thing is restricting you from using the cosine

OpenStudy (trojanpoem):

yea, approximately.

OpenStudy (anonymous):

okay thank you there's your medal

OpenStudy (trojanpoem):

No problem.